1. You have received an order for a dose of 240 mg of theophylline.

What amount of theophylline elixir 80 mg/15 mL is required for the ordered dose?
1
Expert's answer
2015-09-07T13:03:38-0400
If a dose of 80 mg/15 mL theophylline elixir is 240 mg, then 240 mg/(80 mg/15 mL)=(240/80)*15 mL=3*15 mL=45 mL..